Solve the inequality
4k^3-6k^3(< or =)4k
possible answers:
(-infinite,-1/2] or (0,2)
(-infinite,-1/2) or (0,2)
(-infinite,-1/2] or [0,2]
(-infinite,-1/2) or [0,2]

Answer by Theo(13342)

simplify the left side of the equation to get
-2k^3 <= 4k
divide both sides by k to get
-2k^2 <= 4
divide both sides of the equation by -2 to get
k^2 >= -2
take the square root of both sides of the equation to get
k >= square root of (-2)
to prove your answer is correct you need to.
1. let k = square root (-2) and solve. the equation should hold true.
2. let k be > than square root (-2) and solve. the equation should hold true.
3. let k be < than square root (-2) and solve. the equation should be false.